# David Roodman Open Philanthropy San Francisco, CA

### v3

@@ 3.2 Wild bootstrapping
And the functionality in boottest, so revised, has been transplanted to the Julia package WildBootTests.jl. Adding the julia option to a boottest command line switches to the new back end. [^9] Jan Ditzen suggested this interface for calling the Julia implementation.
Section 8.4 of Roodman et al. (2019) approximately replicates two 2SLS regressions in the Levitt (1996) study of the short-term impact of decarceration on crime. [^10] Data and preparation code are available at davidroodman.com/david/Levitt.zip. Here is a version of that example that substitutes reghdfejl for ivregress and sets the stage for calling boottest to perform inference about changes in the size of the prisoner population affect crime rates:
@@
On a single CPU core, version 2.3.5 of boottest, the one published with Roodman et al. (2019), needs 210 seconds to perform the two boottest calls. The latest version finishes in 7.4 seconds, even without calling Julia. Appending julia to the boottest line above lowers the time to 4.1 seconds. (However, the first time boottest is called after installation, there is a significant delay for pre-compilation, just as when jl is first run, as described in section section 2.1.)
@@ 3.3 Plotting
Makie works with several plotting back ends, which render output in formats designed for different applications, such as interactive graphs in web pages and publication-quality image files. The Cairo back end is best for the latter. It is bundled with Makie in CairoMakie.jl:
  • jl AddPkg CairoMakie  # one-time installation
    jl: using CairoMakie  # use now
In 2020, Chuck Huber of StataCorp blogged about using Python to visualize results from a Stata regression.111111blog.stata.com/2020/09/14/stata-python-integration-part-5-three-dimensional-surface-plots-of-marginal-predictions In Huber’s example, a binary indicator of high blood pressure is regressed on age, weight, and their product using the logistic command. margins is then called to compute the probability of high blood pressure according to the fitted model, across a grid of ages and weights. The probabilities are copied to Python and depicted in a three-dimensional surface plot. An adaptation of that example to Julia:
@@
This time I format the Julia commands not as they would be written in a do file, but as they would be typed—or pasted—into Stata’s Command window, That is, I drop the “;” and “///” from the ends of lines. The first Julia line imports the data as a matrix, converts it to a DataFrame, and names the columns. Then the surface plot is built and stored in the Julia variable f. The call to surface() refers to columns of the DataFrame with the convenient dot syntax, as in df.age; here, column names are not prefixed with a colon. The next command displays the figure in a new window, which might pop up behind the Stata window. That line takes advantage of Julia’s piping syntax for function calls, using the symbol “|>”. (display(f) would be equivalent.) Last, the image is saved to disk. It appears in Figure 1.
@@ 4.2 Multithreading
Julia includes several facilities for multitasking. One is distributed processing, which involves launching multiple copies of Julia that can pass messages and data to each other. reghdfejl’s vce(bs) option uses that feature. A lighter form is multithreading, in which multiple execution lines run within one Julia process, largely sharing the same data space. That is the sort @tturbo exploits. The maximum number of threads Julia allows is set when it launches, and it can default to 1. A good guess for the optimum is the number of “performance” or “big” cores on the CPU. You can determine the current limit with the command jl:Threads.nthreads().
How best to set Julia’s thread limit depends on the operating system. In Windows, you can open the Environment Variables control panel and assign JULIA_NUM_THREADS. In Linux, you can add an entry such as export JULIA_NUM_THREADS = 6 to $\sim$/.bashrc” or the equivalent. In macOS, environment variables do not affect graphical applications such as Stata when launched from the desktop. One can instead set the limit in each Stata session through the threads() option of the jl start command. Example: jl start, threads(8). This must be the first jl command issued after starting Stata. It must also precede invocation of Julia-dependent programs such as reghdfejl. [^12] In macOS, one can in fact assign JULIA_NUM_THREADS in $\sim$/.zshenv”. This will take effect when one launches Stata from the terminal. For example open -a StataMP will launch Stata/MP.
@@ 4.3 Limitations in accessing Stata objects
The julia package lets you transfer information between Stata and Julia with commands issued in either language; section 5 will lay out more details. However initiated, the job of copying falls to the C++ plugin included in the julia package. To read and write Stata macros, scalars, matrices, and variables, the plugin must in turn call the official Stata plugin interface. [^13] See stata.com/plugins. The SPI imposes certain constraints. For one, it makes public no functions that can create Stata variables or matrices. As a result, you can edit existing Stata variables and matrices from Julia, but not create new ones. The Stata-side jl subcommands can and do evade this constraint because they start from the Stata environment. For instance, jl GetMatFromMat creates a destination Stata matrix before calling the plugin to copy into it from a Julia matrix.
@@
However, the same complication prevents one from reading Stata locals from Julia. The workaround just described only goes in one direction, because I know of no way for an ado program to read its caller’s local macros.
@@
This trick will not work in an interactive Julia session. [^14] The single, compound line jl: st_local("a","123"); print("‘a’") will also fail. The second command’s reference to a will be expanded before control is passed to Julia. The commands will work as intended if split into two lines.
@@ 4.4 Speed considerations in copying data: type conversions, missingness, and subsample restrictions
StataCorp’s definition of the Stata plugin interface introduces several wrinkles into the copying of data between Stata and Julia.
One is that the functions for reading and writing numeric Stata variables, SF_vdata() and SF_vstore(), only work with double-precision values, regardless of whether a Stata-side variable is stored as byte, int, long, float, or double. By default, the jl commands that copy from the active Stata data set to Julia DataFrames—jl save and jl PutVarsToDF—convert the double-precision values back to the original data types. [^15] Because the ranges of Stata’s byte, int, and long types are narrowed in the way described in the next paragraph, when transferring data from Julia to Stata, the corresponding Julia types are mapped to larger types in Stata. For example, columns of type Int8, which have an allowed range of $[-128,+127]$, are copied to Stata’s two-byte int type, since Stata bytes only hold $[-128,100]$. So a Stata byte variable will be converted to double and back on the way to appearing in a Julia DataFrame. The jl commands can do nothing about the first conversion since that is performed by Stata. But they do accept a doubleonly option to skip the second conversion and instead leave all copied values in double precision. This typically saves time, both by reducing conversions and by allowing Julia to allocate a contiguous block of memory to receive purely double-precision values, rather than a separate block for each of variously typed columns.
Separately, Stata and Julia represent missing values in different ways. Stata reserves certain high values of each data type for 27 flavors of missing: . and .a, …, .z. That is why the maximum value of a Stata byte is 100, not than 127. Julia only supports one flavor of missing, and stores missingness information alongside rather than in data values. It thus distinguishes between empty and missing strings. The jl subcommands translate missings when copying between Stata data sets and Julia DataFrames (but not Julia matrices). For data known to contain no missing values, adding the nomissing option will save time by skipping this step.
Finally, the SPI lets plugin calls include if or in clauses to restrict the sample on the Stata side. When such clauses are not specified, the julia plugin runs specially optimized code.
Data transfers are thus fastest when they map double precision to double precision and include no missing values or sample restrictions. Consider this example, which creates a 10 million $\times$ 10 set of double-precision numbers and copies them to a Julia DataFrame named demo:
@@ 4.5 Simulated read-eval-print loop
The Julia language system has many components—for compiling code, managing memory, and so on. The core parts are written in C++ and distributed as compiled libraries. Other components are themselves written in Julia, including the front end, which users often think of as Julia per se. The front end is called the REPL, for “Read-Eval-Print Loop.” Indeed, the REPL’s core job is to read in commands, evaluate them, print the results, and repeat. But the REPL provides other conveniences. Typing ”?” starts help mode. Typing ”]” instead switches to the Julia package manager. Greek letters and other symbols can be entered through a TeX-like syntax. Arrow keys allow one to explore one’s command history.
@@
And the processing required for the simulation, along with the special handling of Stata locals, adds time to the execution of each Julia line from Stata. The extra 0.02 seconds or so is not noticeable in an interactive session. But it can add up in a program that calls jl: many times. The _jl: variant of the prefix command bypasses all of this REPL simulation.
@@ 5.1 Architecture
The julia package consists of three intertwined pieces. The core is a shared library written in C++ according to the specification for Stata plugins. It comes precompiled for Windows, Linux, and macOS with Intel or Apple silicon. The “julia.pkg” manifest file instructs Stata to select the copy of the library that will run on the target computer, and to name it “jl.plugin.” The plugin performs two main jobs: efficiently copying data among matrices and data sets in Julia and Stata, and sending user-written commands to Julia and returning the results.
@@ 5.2 Stata commands in the julia package
jl also accepts subcommands. As discussed in section 4.2, one is especially useful in macOS, because it allows users to cap the number of threads:
@@
Several subcommands help manage Julia packages, which are discrete libraries of code structured for easy sharing and are usually hosted on GitHub. Just as Stata downloads packages into folders such as “/home/droodman/ado/plus,” Julia downloads and stores packages in directories called environments. Julia also makes it easy to change the current environment. This is useful for the developer of a Julia-calling Stata program because it allows the Stata program to install needed Julia packages without interfering with the user’s package configuration. The package management subcommands are:
@@
[…]
@@
      [a93c6f00] DataFrames v1.6.1
The output ends with a list of currently installed packages—here, the two that are automatically installed.
@@ 5.3 Julia functions in the julia package
Functions in the stataplugininterface module can also be characterized as low- or high-level. The low-level ones wrap and mimic functions in the Stata plugin interface. For example, SF_vdata() extracts a single observation on a single variable. The higher-level ones build on the low-level ones in order to imitate some of Mata’s Stata interface functions. st_view(), for instance, returns a Julia view onto one or more numeric Stata variables, optionally restricting to a subsample. Through the view, Julia code can read and write Stata variables while treating them as part of a matrix.
The full list of Stata-interfacing Julia functions:
@@ 6 Conclusions
Julia is a strong option for universal back-end development of statistical applications. Working in Julia comes with some drawbacks, including the challenge of learning a complex language, the delays upon installation and first use of programs, and the immaturity of some packages and their documentation. But Julia and nearly all Julia packages are free and open and run on all popular operating systems. The language blends features for maximizing performance where it matters with a Python-like expressiveness that makes code easier to write and maintain. The package ecosystem covers domains essential to efficient statistical work, such as multiprocessing, linear algebra, and numerical optimization.
@@
For Stata users and developers, the julia package demonstrates the practicality and value of bridging from Stata to Julia. Through a plugin, data can be piped at high speed. Most features needed to make Stata objects accessible in Julia can and have been implemented. The linkage gives Stata users access to Julia packages for estimation and plotting that offer greater speed or new features. reghdfejl might be the “killer app” for the Julia-Stata link. It fits linear models much faster than the pioneering reghdfe, at least for problems that are time consuming for the older program. It also handily outpaces the newly augmented areg. And, preliminarily, reghdfejl handles nonlinear HDFE models, thus bringing new functionality to Stata users.
StataCorp could strengthen Stata’s bond to Julia, as it has done for Python and Java. Stata’s developers could, for example, extend the plugin interface to allow creation of Stata matrices and variables. They could add a Julia mode to the do file processor, so that Julia code could be embedded without awkward semicolons and line continuation symbols. They might be able to grant more access to Stata variable storage en bloc, so that Julia code could read and write Stata data without copying it one observation at a time, or even without copying it at all. They could give Mata the ability to call Julia functions.

### v4

@@ 3.2 Wild bootstrapping
And the functionality in boottest, so revised, has been transplanted to Julia, in the package WildBootTests.jl. Adding the julia option to a boottest command line switches to the new back end. [^7] Jan Ditzen suggested this interface for calling the Julia implementation.
Section 8.4 of Roodman et al. (2019) approximately replicates two 2SLS regressions in the Levitt (1996) study of the short-term impact of decarceration on crime. [^8] Data and preparation code are available at davidroodman.com/david/Levitt.zip. Here is a version of that example, which substitutes reghdfejl for ivregress and sets the stage for calling boottest. The purpose remains to perform inference about how changes in the size of the prisoner population affect crime rates:
@@
On a single CPU core the version of boottest published with Roodman et al. (2019) (version 2.3.5), needs 210 seconds to perform the two boottest calls. The latest version finishes in 7.4 seconds, even without calling Julia. Appending julia to the boottest line above lowers the time to 4.1 seconds.
@@ 3.3 Plotting
Makie works with several plotting back ends, which render output in formats suited to different applications, such as interactive graphs in web pages and publication-quality images. The Cairo back end is best for the latter. It is bundled with Makie in CairoMakie.jl:
  • jl AddPkg CairoMakie  // one-time installation
    jl: using CairoMakie  // use now
In 2020, Chuck Huber of StataCorp blogged about using Python to visualize results from a Stata regression.999blog.stata.com/2020/09/14/stata-python-integration-part-5-three-dimensional-surface-plots-of-marginal-predictions In Huber’s example, a binary indicator of high blood pressure is regressed on age, weight, and their product using the logistic command. margins is then called to compute the probability of high blood pressure according to the fitted model, across a grid of ages and weights. The probabilities are copied to Python and depicted in a three-dimensional surface plot. An adaptation of that example to Julia:
@@
This time I format the Julia commands not as they would be written in a do file, but as they would be typed—or pasted—into Stata’s Command window. That is, I drop the “;” and “///” from the ends of lines. The first Julia line imports the data as a matrix, converts it to a DataFrame, and names the columns. Then the surface plot is built and stored in the Julia variable f. The call to surface() refers to columns of the DataFrame with the convenient dot syntax, as in df.age; here, column names are not prefixed with a colon. The next command displays the figure in a new window, which might pop up behind the Stata window. That line takes advantage of Julia’s piping syntax for function calls, using the symbol “|>”. (display(f) is equivalent.) Last, the image is saved to disk. It appears in Figure 1.
@@ 4.2 Multithreading
Julia includes several facilities for multitasking. One is distributed processing, which involves launching multiple copies of Julia that can pass messages and data to each other. A lighter form is multithreading, in which multiple execution lines run within one Julia process, largely sharing the same data space. That is the sort @tturbo and the vce(bs) option of reghdfejl exploit. With regard to the latter, the maximum number of threads allowed in each Julia instance is set when the instance launches. It can default to 1. In Stata, you can determine the current limit with the command jl:Threads.nthreads(). A good guess for the optimum is the number of “performance” or “big” cores on the CPU.
How to set Julia’s thread limit depends on the operating system. In Windows, you can open the Environment Variables control panel and assign JULIA_NUM_THREADS. In Linux, you can add an entry such as export JULIA_NUM_THREADS = 6 to $\sim$/.bashrc” or the equivalent. In macOS, environment variables do not affect graphical applications such as Stata when launched from the desktop. One can instead set the limit in each Stata session through the threads() option of the jl start command. Example: jl start, threads(8). This must be the first jl command issued after starting Stata. It must also precede invocation of Julia-dependent programs such as reghdfejl. [^10] In macOS, one can in fact assign JULIA_NUM_THREADS in $\sim$/.zshenv”. This will take effect when one launches Stata from the terminal. For example open -a StataMP will launch Stata/MP.
@@ 4.3 Limitations in accessing Stata objects
The julia package lets you transfer information between Stata and Julia with commands issued in either language; section 5 will lay out more details. However initiated, the job of copying falls to the C++ plugin included in the julia package. To read and write Stata macros, scalars, matrices, and variables, the plugin must in turn call the official Stata plugin interface. [^11] See stata.com/plugins. The SPI imposes certain constraints. For one, it makes public no functions that can create Stata variables or matrices. As a result, you can edit existing Stata variables and matrices from Julia, but not create new ones. The Stata-side jl subcommands can and do evade this constraint because they start from the Stata environment. For instance, jl GetMatFromMat creates a destination Stata matrix before calling the plugin to copy into it from Julia.
@@
However, the same complication prevents one from reading your Stata locals from Julia. The workaround just described only goes in one direction, because I know of no way for an ado program to read its caller’s local macros.
@@
This trick does not work in interactive Julia sessions (ones started by typing jl by itself). [^12] The single, compound line jl: st_local("a","123"); print("‘a’") will also fail. The second command’s reference to a will be expanded before control is passed to Julia. The commands will work as intended if split into two lines.
@@ 4.4 Speed considerations in copying data: type conversions, missingness, and subsample restrictions
Certain details of the Stata plugin interface introduce wrinkles into the copying of data between Stata and Julia.
One is that the functions for reading and writing numeric Stata variables, SF_vdata() and SF_vstore(), only truck in double-precision values, regardless of whether a Stata-side variable is stored as byte, int, long, float, or double. By default, the jl commands that copy from the active Stata data set to Julia DataFrames—jl save and jl PutVarsToDF—convert the double-precision values back to the original data types. [^13] Because the ranges of Stata’s byte, int, and long types are narrowed in the way described in the next paragraph, when transferring data from Julia to Stata, the corresponding Julia types are mapped to larger types in Stata. For example, columns of type Int8, which have an allowed range of $[-128,+127]$, are copied to Stata’s two-byte int type, since Stata bytes only hold $[-128,100]$. So a Stata byte variable will be converted to double and back on the way to appearing in a Julia DataFrame. The jl commands can do nothing about the first conversion since that is performed by Stata. But they do accept a doubleonly option to skip the second conversion and instead leave all copied values in double precision. This typically saves time, by reducing conversions and allowing Julia to allocate a contiguous block of memory to receive purely double-precision values, rather than a separate block for each of variously typed columns.
Separately, Stata and Julia represent missing values in different ways. Stata reserves certain high values of each data type for 27 flavors of missing: . and .a, …, .z. That is why the maximum value of a Stata byte is 100, not than 127. Julia only supports one flavor of missing, and stores missingness information alongside rather than in data values. It thus distinguishes between empty and missing strings. The jl subcommands translate missings when copying between Stata data sets and Julia DataFrames (but not Julia matrices). For data known to contain no missing values, adding the nomissing option to the copying commands will save time by skipping this step.
Finally, the SPI lets plugin calls include if or in clauses to restrict the Stata sample on which the commands will operate. When such clauses are not specified, the julia plugin runs specially optimized code.
Data transfers are thus fastest when they map double precision to double precision and include no missing values or sample restrictions. Consider this example, which creates a 100 million $\times$ 10 set of double-precision numbers and copies them to a Julia DataFrame named demo:
@@ 4.5 Simulated read-eval-print loop
The Julia language system has many components—for compiling code, managing memory, and so on. The core parts are written in C++ and distributed as compiled libraries. Other components are themselves written in Julia, including the front end, which users often think of as Julia per se. The front end is called the REPL, for “Read-Eval-Print Loop.” As its name makes plain, the REPL’s main job is to read in commands, evaluate them, print the results, and repeat. But the REPL provides other conveniences. Typing ”?” starts help mode. Typing ”]” instead switches to the Julia package manager. Greek letters and other symbols can be entered through a TeX-like syntax. Arrow keys allow one to explore one’s command history.
@@
And the processing required for the simulation, along with the special handling of Stata locals, adds time to the execution of each Julia line from Stata. The extra 0.02 seconds or so is not noticeable in an interactive session. But it can add up in a program that calls jl: many times. The _jl: variant of the prefix command saves time by bypassing all of this REPL simulation.
@@ 5.1 Architecture
The julia package consists of three intertwined pieces. The core is a shared library written in C++ according to the specification for Stata plugins. It comes precompiled for Windows, Linux, and macOS with Intel or Apple silicon. At installation, the “julia.pkg” manifest file instructs Stata to select the copy of the library that will run on the target computer, and to name it “jl.plugin.” The plugin performs two main jobs: efficiently copying data among matrices and data sets in Julia and Stata, and sending user-written commands to Julia and returning the results.
@@ 5.2 Stata commands in the julia package
jl also accepts subcommands. As discussed in section 4.2, one is especially useful in macOS, because it allows users to set the number of available threads:
@@
Several subcommands help manage Julia packages, which are discrete libraries of code structured for easy sharing and are usually hosted on GitHub. Just as Stata downloads packages into folders such as “/home/droodman/ado/plus,” Julia downloads and stores packages in directories called environments. Julia also lets you create new environments and switch among them. This is useful for the developer of a Julia-calling Stata program because it allows the Stata program to install needed Julia packages without interfering with the user’s package configuration. The package management subcommands are:
@@
    Current environment: myenv, at /Users/davidroodman/.julia/environments/v1.11/myenv
    Status ´~/.julia/environments/v1.11/myenv/Project.toml´
@@
    [a93c6f00] DataFrames v1.7.0
@@ 5.3 Julia functions in the julia package
Functions in the stataplugininterface.jl module can also be characterized as low- or high-level. The low-level ones wrap and mimic functions in the Stata plugin interface. For example, SF_vdata() extracts a single observation on a single Stata variable. The higher-level ones build on the low-level ones in order to imitate some of Mata’s Stata interface functions. st_view(), for instance, returns a Julia view onto one or more numeric Stata variables, optionally restricting to a subsample. Through the view, Julia code can read and write Stata variables while treating them as part of a matrix.
The full list of Stata-interfacing Julia functions follows:
@@ 6 Conclusions
Julia is a strong option for universal back-end development of statistical applications. True, working in Julia comes with drawbacks, including the challenge of learning a complex language, the delays upon installation and first use of programs, and the immaturity of some packages and their documentation. But Julia and nearly all Julia packages are free and open and run on all popular platforms. The language blends features for maximizing performance where it matters with a Python-like expressiveness that makes code easier to write and maintain. The package ecosystem covers domains essential to efficient statistical work, such as multiprocessing, linear algebra, and numerical optimization.
@@
For Stata users and developers, the julia package demonstrates the practicality and value of bridging from Stata to Julia. Through a plugin, data can be piped at high speed. Most features needed to make Stata objects accessible in Julia can and have been implemented. The linkage gives Stata users access to Julia packages for estimation and plotting that offer greater speed or new features. reghdfejl might be the “killer app” for the Julia-Stata link. It fits linear models much faster than the pioneering reghdfe, at least for problems that are time consuming for the latter. It also outpaces the newly augmented areg. And, preliminarily, reghdfejl handles nonlinear HDFE models, thus bringing new functionality to Stata users.
StataCorp could tighten Stata’s integration with Julia, as it has done with Python and Java. Stata’s developers could, for example, extend the plugin interface to allow creation of Stata matrices and variables. They could add a Julia mode to the do file processor, so that Julia code could be embedded without awkward semicolons and line continuation symbols. They might be able to grant more access to Stata variable storage en bloc, so that Julia code could read and write Stata data without copying it one observation at a time, or even without copying it at all. They could give Mata the ability to call Julia functions.